### v1

# Polarization of reflected X-ray emission from Sgr A molecular complex: multiple flares, multiple sources?

Polarization of reflected X-ray emission from
Sgr A molecular complex: multiple flares, multiple sources?
Extended X-ray emission observed in the direction of several molecular clouds in the Central Molecular Zone (CMZ) of our Galaxy exhibits spectral and temporal properties consistent with the “X-ray echo” scenario. It postulates that the observed signal is a light-travel-time delayed reflection of a short (1.5 yr) and bright () flare, most probably produced a few hundred years ago by Sgr A*. This scenario also predicts a distinct polarization signature for the reflected X-ray continuum, with the polarization vector being perpendicular to the direction towards the primary source and polarization degree (PD) being determined by the scattering angle. We report the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the Imaging X-ray Polarimetry Explorer (IXPE) in 2022 and 2023. We confirm the previous polarization measurement for a large region encompassing Sgr A complex with higher significance, but also reveal an inconsistent polarization pattern for the brightest reflection region in its center. Specifically, X-ray polarization from this region is almost perpendicular to the expected direction in the case of Sgr A* illumination and shows a smaller degree of polarization compared to the large region. Taken at face value, this could indicate the simultaneous propagation of several illumination fronts throughout the CMZ, with the origin of one of them not being Sgr A*. The primary source could be associated with the Arches stellar cluster or a currently unknown source located in the closer vicinity of the illuminated cloud, potentially lowering the required luminosity of the primary source. Although significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios, a combination of high-resolution imaging and micro-calorimetric spectroscopy offers an additional promising path forward.
1 Introduction
X-ray reflection off molecular clouds in the Central Molecular Zone (CMZ) of our Galaxy offers a unique window into the past activity record of the currently dormant supermassive black hole Sgr A* (Vainshtein & Syunyaev, 1980; Sunyaev et al., 1993; Markevitch et al., 1993; Koyama et al., 1996; Ponti et al., 2013). The light-travel-delay due to propagation from the primary source to the reflector implies that the currently observed reflected X-ray emission is connected to the primary’s source luminosity decades to centuries ago (as first realized and described for nova eruptions by Kapteyn, 1901; Couderc, 1939).
The apparent variability of the observed diffuse emission (Revnivtsev et al., 2004; Muno et al., 2007; Inui et al., 2009; Ponti et al., 2010; Capelli et al., 2012; Clavel et al., 2013; Ryu et al., 2013; Zhang et al., 2015; Krivonos et al., 2017; Chuard et al., 2018; Terrier et al., 2018; Kuznetsova et al., 2022; Khabibullin et al., 2022; Brunker et al., 2025; Stel et al., 2025) can be best explained by the short flare scenario – the duration of the primary energetic event responsible for the clouds’ illumination cannot exceed the observed duration of reflected emission from the smallest sub-structures, the light-crossing time111The ‘scanning’ speed of the illumination front propagation along the line of sight varies from at the back side of the illumination paraboloid to infinity in its front side ( is the speed of light). of which is relatively short (e.g., Churazov et al., 2017a).
Since the observed surface brightness of the reflected emission is determined primarily by the distance between the cloud and the source, the mean density of the molecular gas and flare’s fluence (flux integrated over duration of a flare), one can use density estimates obtained from molecular line emission to obtain a lower limit on the average luminosity of the primary source (corresponding to the upper limit on the flare’s duration for a fixed value of fluence). Such estimates give us a limit that still allows a flare from an X-ray binary at the Eddington level, , lasting for one year as a possible origin for the X-ray echo. On the other hand, for an Eddington level flare of Sgr A*, , the duration of merely an hour would suffice to produce the required X-ray fluence. Although large amplitude flares are regularly observed from Sgr A* in infrared and X-ray monitoring campaigns (Eckart et al., 2004; Dodds-Eden et al., 2009; Ponti et al., 2017), the fluence required to account for the X-ray echo is larger by many orders of magnitude and it is unclear if it could have been produced by the same mechanism (e.g., Ponti et al., 2013; Churazov et al., 2017c).
An important and instrumental prediction of the X-ray reflection scenario is linear polarization of the continuum part of the emission, with the polarization degree determined (in the simplest case of single scattering and unpolarized primary emission) by the cosine of the scattering angle and the polarization angle (PA) being perpendicular to the direction towards the primary emission source (Churazov et al., 2002, 2017a; Marin et al., 2014, 2015; Khabibullin et al., 2020b; Di Gesu et al., 2020; Ferrazzoli et al., 2021). Thus, X-ray polarimetry has been long recognized as a potential way to pin-down the location of the primary source, the relative 3D position of the reflecting cloud, and help break degeneracies in parameters of the spectral reflection model (e.g., relative abundance of iron responsible for production of the brightest fluorescent line at 6.4 keV, Churazov et al., 2017b).
The first polarimeteric observation of the currently brightest (in reflected emission) molecular complex Sgr A (Khabibullin et al., 2022; Brunker et al., 2025; Stel et al., 2025) was performed by the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al., 2022) observatory in 2022, broadly confirming the prediction for a short flare from Sgr A* scenario some 200 yr ago (Marin et al., 2023). Even for this brightest complex and rather long effective exposure time of the observation, the level of statistical and systematic noise enabled a polarization measurement only for the whole complex and detection of the polarization at a 3 level.
In order to improve the statistical significance of this result and enable measurements for individual substructures, a second observation was performed in 2023, doubling the total clean (after filtering time intervals affected by enhanced background level) exposure time at the source position (Churazov et al., 2024).
Here we report the results of the two epochs of deep observations of Sgr A taken with IXPE. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with higher significance, but also reveal a possibly inconsistent polarization pattern for the brightest reflection region in its center. Specifically, X-ray polarization from this bright region is almost perpendicular to the expected direction in the case of Sgr A* illumination and shows a smaller degree of polarization compared to the large region.
The paper is organized as follows: we outline the short flare scenario and expected polarization signal in Sect. 2. The Chandra and IXPE data are described in Sect. 3. In Sect. 4, we demonstrate consistency of the observed emission from the Sgr A region with the X-ray reflection scenario using epoch-to-epoch variations of the diffuse emission. We present spatially-resolved spectro-polarimetric measurements from the combined IXPE dataset in Sect. 5. We discuss the results in Sect. 6, and present the summary and outlook in Sect. 7.
2 Short flare scenario
Let us consider a compact molecular complex that is reflecting X-ray emission from Sgr A*’s flare that occurred yr ago. Except for the photoionization followed by fluorescent line emission (most importantly the fluorescent line of the neutral iron at 6.4 keV), the key process is Compton scattering, which, in the majority of the relevant cases, might be treated in the optically thin limit (e.g., Churazov et al., 2017b). The polarization of the reflected X-ray continuum is fully determined by the geometry of the scattering, set by the relative disposition of the cloud and the primary source, if the primary radiation was unpolarized.
We use the coordinate system such that the primary source is located at the origin, the -axis is pointing from the observer towards it, and the reflecting cloud resides in the -plane (see Fig. 1). This plane is referred to as the scattering plane in what follows, and the -axis represents the direction normal to it.
The line-of-sight and image plane coordinates of the reflecting gas are connected by the relation
which can be formulated in the dimensionless (and time-independent) form after dividing by the characteristic length scale pc , so that :
The distance from the primary source to the cloud equals
and the cosine of the scattering angle (see Fig. 1) is calculated as
For the flare duration yr , the thickness of the gas layer contributing to the reflected emission for any line-of-sight direction is small
so one can neglect variations in the scattering geometry across this layer.
The degree of polarization depends on the cosine of the scattering angle (in the limit of Thomson scattering) as
Thus, knowing the degree of polarization , one can determine the 3D position of an illuminated cloud. For instance, measured for the Sgr A complex by Marin et al. 2023 corresponds to , and for one gets , i.e. the illuminated cloud is further away from us than Sgr A∗ by this distance.
One can explicitly relate the delay time and the measured degree of polarization
This expression will be used to estimate the age of the flare.
An extension of this basic scenario, allowing for possible intrinsic polarization of the original flare’s emission, was considered by Churazov et al. (2002), who presented explicit expressions for the polarization degree of the scattered emission as a function of the cloud’s relative position (see also Gilfanov et al., 1987, where scattering of anisotropic and polarized time-variable radiation has been modeled). This analysis was expanded by Khabibullin et al. (2020b) to predict the full polarization pattern of the scattered emission, i.e., the maps of the polarization degree and polarization plane orientation for various configurations of the primary source polarization. In contrast to the case of the unpolarized primary emission, the predicted patterns demonstrate substantial morphological complexity, including non-axisymmetric variations in the intensity, polarization degree, and polarization plane orientation. For instance, in this case, the electric field vector is not necessarily perpendicular to the line connecting the primary source and the reflecting cloud. Also, given that, contrary to the scattered continuum emission, fluorescent line emission is insensitive to the intrinsic polarization of the initial flare, correlated spatial variations in polarization and spectral properties (e.g., equivalent width of the fluorescent line) are expected in this case (Khabibullin et al., 2020b).
3 Data
High spatial resolution X-ray images were obtained with the Chandra Observatory (Weisskopf et al., 2000) over multiple observing campaigns from 2000-2022. The total exposure of the Chandra pointings with G0.13–0.11 located in the central region of the field of view (FOV) with high spatial resolution is 1.4 Msec. The Chandra data reduction follows a standard procedure based on the latest versions of the data reduction software (CIAO v. 4.14) and calibration (CALDB v. 4.9.8). Our particular approach and analysis steps are described in detail in Vikhlinin et al. (2009).
Briefly, they include the identification and removal of high background periods, the correction of photon energies for the time and detector temperature dependence of the charge transfer inefficiency and gain, and the creation of matching background datasets using blank sky observations with exposure times similar to the GC pointings. For the analysis presented here, we used the combined flat-fielded and background-subtracted Chandra images in the 3–8 keV band and spectra extracted in several regions of interest. Following the standard approach for analyzing Chandra spectra of extended sources, we have generated the spectral response files that combine the position-dependent ACIS calibration with the weights proportional to the observed brightness.
In this work, we used the data of two IXPE observations of the Galactic Center region, more specifically, a complex of molecular clouds degrees to the east of Sgr A* (Marin et al., 2023) – the Sgr A complex. These observations with a total clean (i.e., after excluding periods of enhanced particle background and solar flares) exposure time of Msec, were performed in two parts – in February 2022 and September 2023. These two observations were targeted at (RA, Dec)=(266.51) and (266.57), respectively, with an intentional offset between their aim-points of .
The raw IXPE data were processed with the standard pipeline (version 1.9.2 with HEASOFT v6.34)222https://heasarc.gsfc.nasa.gov/docs/software/heasoft/ and CALDB version of May 2025. The output FITS files of this pipeline contain the event-by-event Stokes parameters (see Kislat et al., 2015) from which the polarization observables of the X-ray radiation can be derived. The data products are publicly available for use by the international astrophysics community at the High-Energy Astrophysics Science Archive Research Center333https://heasarc.gsfc.nasa.gov/docs/ixpe/archive/ (HEASARC, at the NASA Goddard Space Flight Center). An energy-dependent particle background rejection algorithm (IXPE-Background444https://github.com/aledimarco/IXPE-background tool, v2.3) was applied to these level-2 event files; it allows the removal of of the instrumental background (Di Marco et al., 2023). Observation periods affected by increased background due to solar activity were removed (see Marin et al. 2023 for additional details on the procedure). The second epoch observation was performed with a smaller dithering amplitude, aiming at achieving a higher sensitivity in the central part of the field of view. As a result, the brightest compact (i.e., spatially unresolved) source, the Arches cluster, which was used to adjust the astrometry of the first observation, was not visible in the nominal field of view. The Arches cluster was, however, detectable near a corner of the field of view of Detector Unit 1 (DU1) in the Level 1 (L1) event data. Custom analysis of L1 data to determine attitude offsets for the individual observation segments yielded a means for consistent astrometric correction. These attitude corrections were applied after the standard pipeline processing to generate the final Level 2 (L2) products, used here and available in the HEASARC. The resulting astrometry is fully sufficient for robustly singling out the emission of the brightest reflection region, as well as pulsar wind nebula (PWN) G0.13-0.11. No additional weighting or binning of the data was included in the spatial (broad-band) or spectropolarimetric analyses. Long IXPE observations of the Circinus galaxy were used to model quiescent background spectrum, i.e. the sum of the sky and detector backgrounds in the area not contaminated by the target source (after applying identical filtering routines), and subtract it in the spectral analysis of the Stokes data (consistently with the procedures performed and described in Marin et al. 2023).
4 X-ray images, light curves and spectra: epoch-to-epoch variations from Chandra
The short flare reflection scenario is well established and tested on the data of extensive Chandra and XMM-Newton monitoring campaigns (e.g., Ponti et al., 2010; Clavel et al., 2013; Churazov et al., 2017c; Terrier et al., 2018; Ponti et al., 2013). Here, we are interested in the quantitative characterization of the reflection signal from several selected regions in order to use this as a basis for the spectropolarimetric analysis of the IXPE data.
4.1 Imaging
Due to the expected spectral shape of the reflected emission (Sunyaev & Churazov, 1996; Churazov et al., 2017a) and high foreground interstellar absorption towards the Galactic Center, the optimal band is commonly selected as 3–8 keV. This band contains both the reflection continuum and the most prominent fluorescent line of neutral (or only weakly ionized) iron at 6.4 keV. Also, it provides good leverage for spectral de-composition analysis needed to single out the ‘hot plasma’ ( keV) component (bulk of which is produced by unresolved population of accreting white dwarfs, so called Galactic ridge and bulge emission, Revnivtsev et al., 2009), as well an intermediate temperature component (2–3 keV), which is especially bright in the direction of the Sgr A complex. Finally, this band is almost insensitive to the dramatic decrease in Chandra’s response efficiency at low energy due to the build-up of the absorbing contaminant layer, allowing one a simple comparison between the epochs just at the band-integrated maps level (e.g., Churazov et al., 2017c).
The X-ray images of the Sgr A region in the 3–8 keV band obtained by Chandra in different epochs starting from 2013 are shown in Fig. 2. Clear global and local changes in the morphology are visible over this 10-year-long period, even though the whole region stays relatively bright over its entire duration. This behavior is indeed expected, given that 10 years corresponds to the light travel distance of 3 pc (1.27 arcmin at the Galactic Center distance), which is bigger or comparable to the physical sizes of individual substructures ( pc), but smaller than the full complex itself ( pc).
4.2 Light curves
Epoch-to-epoch light curves of the emission in the 3–8 keV band for several regions of this field are shown in Fig. 3. One can see that the average surface brightness of the large circular region encompassing the Sgr A complex (black line) from 2000 to 2024 is relatively stable, even though the morphology of the emission within it was changing. Slight ( level) variations are present, some of which are caused by different positioning of this relatively large region in Chandra’s field of view, resulting in different weighting of various parts of it due to spatially-dependent vignetting and locations of the chip gaps, as well as the contribution of time-variable and transient point sources. This is illustrated by the presence of similar amplitude variation in a background region, which appears to contain a much smaller amount of reflected emission.
The other three regions are smaller in size and show substantial variations in surface brightness, with the brightest central part of the Sgr A complex demonstrating a rapid rise from 2013 to 2015 and then a decay after 2023. Decay of the diffuse emission surrounding the Arches cluster is well known and is also (partially) attributed to the decay of the reflected emission (e.g., Capelli et al., 2011; Krivonos et al., 2014; Clavel et al., 2014; Krivonos et al., 2017; Kuznetsova et al., 2019). More surprising is a decrease in surface brightness from the region dominated by the non-thermal emission from G0.13-0.11 nebula. However, spectral analysis of the variable part of the diffuse emission shows that it is also the decay of the reflection emission that drives this apparent trend (see Sect. 4.3).
4.3 Spectra
Spectroscopy of the time variable emission is a very powerful tool that allows us to get an idea about the spectral shape of the variable and non-variable components, which can then be used for single epoch spectral decomposition. In order to make this method most efficient, we select two epochs of observations that correspond to the minimal and maximal surface brightness in the regions of interest. We focus on three regions well within the Sgr A complex shown in Fig. 4 and the difference between observations taken in 2004 and 2022.
We extract spectra in each of them and fit simultaneously by a model (described and exploited for Sgr A complex in Marin et al., 2023), allowing only the reflection component (derived from CREFL16 model555https://wwwmpa.mpa-garching.mpg.de/~churazov/crefl by Churazov et al., 2017a) to vary. The other components (‘the hot ( keV) plasma’ and ‘warm ( keV) plasma’) are essentially phenomenological descriptions of the temporally constant and spatially smooth emission of background and foreground sources (see the model definitions in Marin et al. 2023). Spectral analysis is performed using XSPEC666https://heasarc.gsfc.nasa.gov/xanadu/xspec/ software package (Arnaud, 1996, V12.14).
Very good quality of the fits (=685 and 705 for 705 degrees of freedom for the ‘bg’ and ‘re’ reflection-dominated regions, and =740 with 705 degrees of freedom for the G0.13-0.11:‘nt’ region due to the presence of a non-thermal powerlaw component in this case) proves that the variable emission has the spectral shape that is indeed fully compatible with the reflection scenario (as shown in Fig. 5). In particular, the difference spectrum contains a hard X-ray continuum and a prominent fluorescent iron line that are treated separately during the fit to allow more flexibility in the spectral shape. The resulting non-reflection emission can be compared with the prediction of the spatial-decomposition model, which is capable of estimating the brightness of the bulge and ridge emission based on the projected stellar mass maps (Anastasopoulou et al., 2025). As a proxy for the spectral shape of this emission, we use spectra extracted from a deep observation of the bulge region (Revnivtsev et al., 2009) which is believed to contain negligible contribution of the diffuse reflected emission (even though featuring the 6.4 keV line due to the presence of such a line in spectra of certain types of accreting white dwarfs, see, e.g., Anastasopoulou et al. 2025). One can see that the emission predicted by this model does not exceed the derived non-variable part of the spectrum for these regions, ensuring that there is no overestimation of the purely reflection component.
We can also see that even for the region dominated by non-thermal powerlaw emission of the G0.13-0.11 nebula, the spectrum of the emission responsible for the drop in brightness since 2004 is fully consistent with the reflection model. Interestingly, the relative strength, or equivalent width, of the fluorescent line differs slightly from region to region. Although this might be indication of a different reflection geometry, it is also within expected variations due to inhomogeneities in iron abundance or relative contribution of second scatterings (it is expected to be larger for fainter regions and feature a relatively stronger fluorescent line with a Compton shoulder profile)777Once the primary photons leave the cloud, it fades, but the doubly and multiply-scattered photons might still be detected with deep observations ( see, e.g. Sunyaev & Churazov, 1996; Churazov et al., 2002; Molaro et al., 2016; Kuznetsova et al., 2022; Khabibullin et al., 2022).. In the case of abundance variations, the predicted polarization signal is unchanged, while the other two cases predict correlated change in the polarization degree ( see, e.g., Churazov et al., 2002; Khabibullin et al., 2020a). In both of them, stronger line emission should correspond to a smaller polarization degree of the continuum, either as a result of contamination by less polarized second scattering emission or due to the phase function of the Compton scattering (for the spectral band of interest here, Thomson limit of the Compton scattering provides a fully adequate description) as a function of the cosine of the scattering angle. Another cause of spatially variable equivalent width of the fluorescent line might be intrinsic polarization of the primary emission, which in principle allows one to produce a high polarization degree in the continuum and a relatively strong emission line, but only at the cost of a relatively low continuum emission overall (Khabibullin et al., 2020a). In this case, the direction of the polarization vector might be very different from the unpolarized primary source prediction.
Imaging, timing, and spectral analysis of the Chandra data allowed us to quantify reflected continuum properties and predict the polarization signal we could expect as a function of the reflection geometry parameters. We are now in a position to combine this with the IXPE data to derive geometrical constraints based on the spatially resolved spectropolarimetric measurements.
5 Spatially resolved X-ray spectropolarimetry
The angular and spectral resolution of IXPE, in addition to the limited time coverage of the Galactic Center region, is not sufficient to disentangle different emission components contributing to the observed signal in an independent manner. We have to rely on spectral and spatial information derived from quasi-simultaneous Chandra observations combined with the historical Chandra and XMM-Newton data, setting the baseline for the non-variable diffuse emission components. In this regard, the first step needed to be done is checking the consistency of the total (i.e., not component-separated) imaging and spectral properties in IXPE Stokes and Chandra data.
5.1 IXPE Stokes images and spectra
Figure 6 shows a comparison of Chandra and IXPE 3–8 keV images for the region of interest. The Chandra image corresponds to the middle 2022 epoch, while the IXPE image combines both 2022 and 2023 observation epochs. Only small morphological variations between 2022 and 2023 observations are expected from a comparison of the Chandra images taken in these epochs. The consistency of Chandra and IXPE images is very good (see Fig. 6), both in the large-scale diffuse emission and the three brightest individual regions – the Arches cluster, G0.13-0.11 PWN, and the brightest reflection spot in Sgr A complex. Some differences in morphology or relative brightness of certain structures are expected, given that the energy dependencies of the effective area of the two telescopes are different, resulting in slightly different weighting of various emission components in the combined 3–8 keV image. In what follows, we consider the combined IXPE dataset and demonstrate consistency of differences between the epochs with the expected statistical deviation level in Appendix A.
The consistency of the IXPE Stokes and Chandra data can be further validated by comparing spectra for several regions of interest. Figure 7 shows the IXPE Stokes spectra for two mutually exclusive regions: a large (r=4.5) circle region encompassing Sgr A (also excluding the region around the G0.13-0.11 nebula) and the brightest reflection spot in its center. Both spectra are perfectly fit with the models derived from the quasi-simultaneous Chandra data with only a slight normalization correction (factor 0.85) needed, which is well within the possible margin of the effect introduced by proper response function calculations for extended sources with complicated morphological structure. This factor is propagated further in a self-consistent manner, so the polarization degree calculations are not affected by this.
Figure 7 also shows, separately, contributions of the reflected (dashed lines) and non-reflected (dotted lines) emission. It is the continuum part of the reflection continuum that is expected to be polarized, and these reflection continuum models are used in the spectropolarimetric analysis later on. One can see that the reflection continuum component of the brightest region is 3.55 times brighter for the central region compared to the average over the large circle. Since the fluorescent line emission of the reflection component should not be polarized, we focus the polarimetric analysis on the 3–6 keV band, which maximizes signal-to-noise ratio for the reflection continuum. We summarize parameters of the reflection models for these regions and their relative contributions to the 3–6 keV band in Table 1.
| Region | Normcont | Normfluor | Frac3-6,cont | ||||||
| per arcmin2 | per arcmin2 | ||||||||
| Sgr A, (excl. br&nt) | 7.0 | -0.3 | 1.0 | 0.24 | |||||
| Sgr A, bright | 0.46 |
5.2 IXPE Stokes and images and spectra
Given the consistency of the morphological and spectral properties of the total intensity (Stokes ) of the diffuse emission measured by IXPE, we are in a position to proceed with mapping and spectral analysis of the polarization signal itself, which is most conveniently expressed in terms of the Stokes and parameters. These are inferred for each individual photon (event), detected by IXPE, and the expectation value for an unpolarized source (point-like or diffuse) is zero (in the absence of systematic errors).
Linearity of the Stokes parameters allows one to produce maps in (optimal) energy bands and smooth or bin them in order to suppress statistical noise. In Fig. 8, we show Stokes and images obtained from the full IXPE data set in the 3–6 keV band, where the signal-to-noise ratio for the reflected continuum is expected to be the highest. One can see dominance of the positive and for the G0.13-0.11 region, close to zero and negative for the large Sgr A region, and marginally positive and for the brightest reflection region. Also, the level of fluctuations due to background noise (and its increasing trend towards the boundaries) is visible. The latter effect is well known and described, arising from imperfections of the photoelectron tracks reconstruction near the boundaries of the Field-of-View (e.g., Di Marco et al., 2023). Our choice of the extraction regions ensures that the possible impact of this effect is minimal, given that all regions are a) always located close to the centers of the detectors; b) are large enough to fully encompass diffuse structures which are bright in reflected emission (even taking into account effective PSF of IXPE). We try to quantify the possible amplitude of other systematic effects in the Appendix A.
The described picture is consistent with the previous findings for the large Sgr A region and the non-thermal emission from G0.13-0.11, but the result obtained for the brightest reflection region is very surprising. Indeed, the Stokes and pattern for it appears more similar to the G0.13-0.11 region than to the Sgr A complex, meaning the direction of polarization is almost perpendicular to the expected one.
We proceed with the spectropolarimetric analysis in order to quantify and explore this in more detail. Following the approach used in Marin et al. (2023) and Churazov et al. (2024), we construct energy spectra of the Stokes values and extracted from the regions of interest and supplemented with corresponding response matrices. In addition to the effective area of the mirror system and efficiency of the detector system, the response matrices take into account the energy-dependent modulation factor, i.e., efficiency of the response to a linearly polarized signal.
Since we expect only the continuum part of the reflected emission component to be polarized, we build a model (Marin et al., 2023; Churazov et al., 2024) that allows one to simultaneously fit Stokes and data, varying polarization degree (PD) and polarization direction (in terms of the polarization angle ) for a fixed spectral shape and normalization of the reflected continuum for each region (found consistently in Chandra and IXPE Stokes spectral analysis described in Sect.s 4.3 and 5.1). The fitting is done via a standard steppar procedure in in XSPEC which involves calculating statistics on an extensive grid of PD and values ranging from 0 to 100% (defined with respect to the spectral model of the reflection continuum) and from to 90 degrees (in IAU convention), respectively. Namely, the model is composed of two additive components with tied spectral shape (given by the spectrum of the reflected continuum) and normalization (set by PD) multiplied by (for Stokes ) and (for Stokes ) user-defined functions. The results of this procedure are shown in Fig. 9. As already hinted by the maps of the Stokes parameters Q&U, the signal from the large Sgr A region is characterized by negative and close to zero values, while for the brightest reflection spot, both and are marginally positive with large statistical noise.
5.3 Inferred polarization parameters
The inferred confidence regions for the large circle (excluding the G0.13-0.11 and the brightest central region) and the brightest central region are shown in Fig. 10 along with the expectations for the reflection scenario with Sgr A* being the primary source (which sets PA, i.e. in this plot) and a range of the scattering angles (which sets the PD). The contours mark levels of 2.3, 4.61, 9.21 with respect to the best-fit parameters corresponding to 68%, 90%, and 99% confidence levels for two fitted parameters.
The best fit values for the large region are and deg (quoted uncertainties correspond to confidence interval), fully consistent with the first measurement reported in Marin et al. (2023), who found and deg. The detection of non-zero polarization is at the level of 3.9 significance based on the combined dataset. A slight increase in the PD is associated with the exclusion of the brightest reflection region, which appears to have almost orthogonal polarization direction (if any, see below), even though the polarization parameters inferred for the large region with and without exclusion of the brightest reflection spot are consistent within the uncertainties. Since this exclusion is made based on intensity only, without any prior knowledge of its polarization properties, no intentional bias is expected to be introduced by this procedure.
Our results for the large region signify the direction to the primary source, fully compatible with the Sgr A* scenario, and the age of the flare is now constrained to yr in this case (corresponding to the cosine of the scattering angle (the solution branch corresponding to positive would correspond to the age of yr, which is unlikely given historically available observations of the Galactic Center region in X-rays).
As already hinted by the Stokes maps, the polarization signal extracted from the brightest reflection region is rather different: the best fit value for the PA is deg and polarization degree , i.e., significance of the detection of non-zero polarization falls just short of the two sigma level. On the other hand, the same polarization parameters as inferred for the large Sgr A region can be excluded at more than the 99% confidence level, given that the best fit polarization directions are almost perpendicular to each other. The 99% upper limit on the PD for the PA fixed at the direction towards Sgr A* is (1 upper limit is ), meaning that a situation with Sgr A* being the primary source and the cloud being significantly more distant (larger ) is still possible.
Given that the brightest reflection region appears to be encompassed by more extended diffuse emission with different polarization, one may assume that there is a contribution of the large-scale emission to the extracted signal from the brightest reflection region. In such a case, one would need to subtract (or model) this contribution when fitting the polarization spectra. Using the signal from the large region as an estimate for the background for the brightest reflection region, the fitting procedure results in (i.e., 2.4) and deg. Naturally, this increase in significance is driven partially by the fact that polarization in the background region is almost perpendicular to the marginally detected polarization in the brightest reflection region. The 99% upper limit on the PD for the PA fixed at the direction towards Sgr A* is (1 upper limit is ) in this case.
Finally, one might expect the impact of systematic effects for the brightest region in the center of the Field-of-View. Analysis of the individual datasets (two observational epochs and 3 DUs) does not reveal exceptional outliers, although a significant detection is achieved only in the combined data for one of the detector units (DU1). For the other two detectors, the hypothesis of polarization signal being identical to the one measured from the large region is excluded at the confidence level of at least 63% and 74% confidence levels (see Table 3).
6 Discussion
The main driver of the long IXPE observations was to map the polarization signal from individual spatially resolved structures, which are bright in reflected emission. We (marginally) succeeded with this, but at the price of revealing a more sophisticated picture, complicating both data analysis and interpretation. Figure 11 presents a summary of the polarization geometry inferred for the three main regions in the Sgr A direction.
While confirming that the total polarization signal from the large region centred on the Sgr A complex might come from Sgr A*, we found a significantly different polarization signal for the brightest and highly time-variable central part of it.
We consider two options that, within uncertainties, are consistent with the data. Option 1: The signal from the compact bright region has the same polarization direction as in the large region, but a significantly lower PD. In this case, accumulated statistics still allow significant fluctuations of the PA, even though the probability of this is low. Option 2: The PA is genuinely different from that inferred for the large region.
The former case leaves open a possibility of Sgr A* being the primary source of illumination, but implies that the brightest reflection spot is located at another distance (lower PD implies larger distance along the line of sight), lending further support to the hypothesis of several flaring episodes (as already suggested in Clavel et al., 2013, based on the spatial and temporal variability patterns).
The latter case can be reconciled with a Sgr A* flare only if its primary emission was strongly polarized with the sky projection of the electric field vector aligned with the source-cloud direction (as shown in the bottom panel of Figure 2 and in Figure 4 of Khabibullin et al., 2020a). This configuration is possible only when the cloud is located at the far side of the reflecting paraboloid, allowing it to ‘mirror’ the intrinsic polarization of the primary source. Although infrared emission of the regular Sgr A* flares is known to be polarized (Eckart et al., 2006; Shahzamanian et al., 2015), and the spatially resolved emission from the currently present accretion flow has been measured (e.g., Event Horizon Telescope Collaboration et al., 2024), the possibility of extrapolating these measurements to a hypothetical orders-of-magnitude brighter accretion episode (e.g., Dovčiak et al., 2008), which might be responsible for the observed X-ray echo, remains uncertain.
At the same time, the spectral analysis of the variable diffuse emission from both large and the brightest regions indicates that the spectral shapes of the reflected emission are rather similar, with the slight variations in the equivalent width of the fluorescent line of neutral iron well within the possible effects of different scattering angle or metallicity of the reflecting cloud. This appears to disfavor intrinsic polarization of the primary emission since, in this case, apparent rotation of the polarization vector would be associated with substantial variations in the ratio of the fluorescent line to the reflected continuum. In other words, the similarity of the observed spectral shape indicates that the scattering angle in both cases should be similar as well. Taken at face value, this would indicate the presence of at least two primary sources responsible for the current X-ray echo(es) propagating through the Sgr A complex. While for the large region Sgr A* remains a very viable candidate, the polarization vector of the brightest region is perpendicular to the direction towards the Arches stellar cluster, which is one of the youngest massive clusters in the Milky Way, and is known to host highly variable compact sources.
The historical reflection emission from Sgr B2, Sgr C, and Sgr D complexes (Terrier et al., 2018; Chuard et al., 2018) is most naturally explained by a flare from a powerful source in the center of the CMZ, i.e., Sgr A* (e.g., Ponti et al., 2013; Churazov et al., 2017a). At the same time, the decay of the reflected emission from these clouds, on a time scale comparable to their light-crossing time, indicates that their illumination was also by a relatively short flare, lasting less than 10 years or so. Also, the spatial distribution of the reflection signal, averaged over 20 yr of available high resolution observations with Chandra and XMM-Newton, shows that not all molecular complexes in the CMZ were illuminated during this period of time, meaning that the coverage of the CMZ by the illumination front is incomplete, indeed corresponding to a relatively thin volume between two paraboloid surfaces containing only some of the largest molecular complexes (Ponti et al., 2013; Churazov et al., 2017a; Terrier et al., 2018; Khabibullin et al., 2022; Anastasopoulou et al., 2025).
Thus, we come to a picture of a plausible ‘global’ illumination of the large molecular complexes in the CMZ by a luminous flare in its centre, and the presence of ‘local’ time–and space-variable reflection signal(s). Although such a picture means a great increase in complexity compared to the simplest ‘single source, single (short) flare’ scenario, it appears to be a possible alternative.
The Galactic Center region is characterized by the highest volume density of all kinds of compact sources, including low-mass and high-mass X-ray binaries and magnetars (e.g., Sazonov et al., 2020). Thanks to the decades-long monitoring of this region by several generations of X-ray observatories, we know that many of them are highly variable, and some are transient. The largest concentrations of such objects are expected to be associated with the massive star clusters in this region, namely the Nuclear Star Cluster, Arches, and Quintuplet clusters. While the former hosts a supermassive black hole Sgr A*, the latter two might host intermediate mass black holes in their centers. A spontaneous accretion event (e.g., due to a partial or total tidal disruption event of a star, planet, or a gas cloud) can, in principle, result in a months-long activation with the luminosity exceeding (e.g., Zubovas et al., 2012).
Contrary to the case of Sgr A* being a primary illumination source, it is more difficult to estimate the required fluence of the initial X-ray flare because of the (more) uncertain relative distance between the primary source and the reflector. In particular, for a source located close to the reflecting cloud, the required fluence of the flare will be even smaller than the value inferred for Sgr B2 or Sgr A complexes in the Sgr A* scenario. This opens the possibility for even less dramatic transient events to be responsible for the observed ‘local’ reflection echoes. Further monitoring of the morphological and spectral changes (e.g., Ponti et al., 2010; Churazov et al., 2017c), in combination with the data on the molecular gas distribution in velocity space (e.g., Ponti et al., 2013; Clavel et al., 2013; Churazov et al., 2017b; Khabibullin et al., 2020a; Brunker et al., 2025), should help us to pinpoint the location of the primary source with more confidence, and, as a result, better quantify required energetics of the flare and its possible nature.
7 Conclusions
We report the results of two deep observations of the currently brightest (in reflected emission) molecular complex Sgr A taken with the imaging X-ray polarimeter IXPE in 2022 and 2023, and supplement them with the data of a quasi-simultaneous Chandra monitoring campaign of the same region. We confirm the previous polarization measurement for a large region encompassing the Sgr A complex with high () significance, but also reveal an inconsistent polarization pattern for the brightest reflection region in its center.
This compact region has a smaller PD compared to the large region, and a PA that is almost perpendicular to the expected direction in the Sgr A* unpolarized-flare scenario. While the statistical significance of this result is at the level of less than 3, it might indicate that several illumination fronts have propagated through the CMZ simultaneously, with at least one of them not due to a Sgr A* flare. The scenario of Sgr A* being the primary source of illumination for both regions is still possible, but requires that either (i) currently the brightest compact region of the Sgr A complex is located at a significantly farther distance (and hence is illuminated by an older flare), or (ii) that the primary emission was itself strongly polarized in a manner allowing almost perpendicular rotation of the polarization vector relative to the unpolarized case (e.g., in the configuration illustrated in Khabibullin et al., 2020a). In fact, the latter option also requires the reflecting medium being well ‘behind’ Sgr A*.
If instead the inferred polarization direction is indeed determined by direction to the primary source, it could be associated with the Arches stellar cluster or a currently unknown transient source located in a closer vicinity to the illuminated cloud (see a discussion of implications of such a scenario in Sunyaev & Churazov, 1998). Although a scenario of multiple (transient) illuminating sources is straightforward to model (e.g., Molaro et al., 2014), the arising complexity of spectral variability and polarization patterns requires sensitive monitoring observations of many individual reflecting complexes (e.g., Marin et al., 2014).
Significantly deeper observations with IXPE would be required to unequivocally distinguish between the scenarios. Given the consistency of the observed noise levels with expected statistical fluctuations, accumulation of an additional 1.1 Ms of clean exposure time would be needed to ensure the confidence level detection for the brightest reflection region (if the real signal is indeed at the detection level and stays constant in time). A combination with continued high-resolution imaging, micro-calorimetric spectroscopy (e.g., with XRISM observatory, Tashiro et al., 2025), and global monitoring of the new reflection regions via CMZ-wide surveys (e.g., Khabibullin et al., 2020a; Semena et al., 2024) offer a promising path forward (along with the roadmap outlined in Churazov et al., 2019). Next-generation X-ray polarimetry missions will be required to disentangle the full complexity of X-ray reflection signals if the scenario of multiple sources and/or multiple flares is confirmed. An additional window might be offered by measuring polarization and spectra at higher energies, ¿10 keV, where the reflected continuum should be dominant over other emission components. Prospective missions such as the Enhanced X-ray Polarimetry Observatory (EXPO), evolution of the NGXP concept (Soffitta et al., 2021), would be ideally suited for this task.
Acknowledgements
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC).
We thank Dylan Maurel of the IXPE Science Operations Center for manually reprocessing the IXPE data.
IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. JS thanks GACR project 21-06825X. AV, RK, and WF acknowledge support from the Smithsonian Institution and the Chandra High Resolution Camera Project through NASA contract NAS8-03060. IL was funded by the European Union ERC-2022-STG – BOOTES – 101076343. RF, ECo, ADM, PSo, FLM, FMu and SF are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”.
Appendix A Systematic uncertainties
The IXPE observations of the Sgr A complex are aimed at faint, spatially and temporarily variable diffuse emission in the vicinity of a number of bright point-like and extended sources. As a result, measurement of the polarization signal and especially mapping of its spatial morphology is very challenging and sensitive to uncertainties in the current knowledge of systematic effects inherent to such observations.
Here we estimate the level of systematic uncertainties that might be present in the data in addition to the purely statistical noise caused by limited count number statistics. In order to do so, we perform a number of jack-knifing experiments by considering variations in the obtained results if various subsets of the full IXPE dataset are included. In particular, one can compare data collected in two epochs separately, by separate Detector Units (DU) of IXPE, as well as spatial variations in excess of the expected statistical noise level. We calculate variance within regions of relevant size (comparable to the brightest in the reflection region) shown in Figures 12 and 13 for Stokes and , respectively.
The signal-to-noise ratios obtained by dividing the extracted average Stokes values from the regions of interest (the brightest reflection spot and G0.13-0.11 region for reference) by the estimated variances from the maps are consistent with the ones obtained from spectral fitting. None of the observation epochs appears to be different from the full dataset in excess of the expected variance, which is estimated by analyzing variances between different halves of the data selected randomly. Thus, we conclude that statistical properties of the Stokes and maps are consistent with expected uncorrelated noise behavior (at least on the spatial scale considered here), meaning that background and signal fluctuations dominate over possible systematic effects.
In order to quantify possible systematic effects of the observational setup or detectors, we also quantify a variance in the polarization signal inferred from the data of individual observation epochs and detector units. Each data subset contains times smaller effective exposure time, boosting the statistical noise in each individual measurement, especially for the Stokes and maps. No significant excess variation of any detector signal in any epoch is observed in this test as well.
| Dataset | , % | , % | PD, % | PA, ° | CL, % | |
| Obs 1 | 1.0±1.4 | 5.7±1.4 | 6.0±1.4 | 50±7 | 4.0 | 99.97 |
| Obs 2 | 1.8±1.4 | 3.4±1.4 | ¡7.5 | NC | 2.7 | 97.47 |
| Combined | 0.4±1.0 | 4.5±1.0 | 4.6±1.0 | 42.5±6 | 4.5 | 99.99 |
| Combined DU1 | 1.1±1.7 | 3.9±1.7 | ¡8.5 | NC | 2.4 | 93.97 |
| Combined DU2 | 1.3±1.8 | 6.2±1.8 | 6.0±2.0 | 39±8 | 3.6 | 99.82 |
| Combined DU3 | 1.3±1.8 | 3.5±1.8 | ¡8.0 | NC | 2.2 | 90.23 |
| Dataset | , % | , % | PD, % | PA,° | CL, % | CL0, % | |
| Obs 1 | 4.3±4.9 | 8.2±4.9 | ¡22 | NC | 1.9 | 83.06 | 98.59 |
| Obs 2 | 2.8±5.1 | 4.9±5.1 | ¡19 | NC | 1.1 | 45.55 | 79.91 |
| Combined | 0.9±3.5 | 6.6±3.5 | ¡16 | NC | 1.9 | 83.18 | 99.05 |
| Combined DU1 | 10.9±6.0 | 16.5±6.0 | 20±6 | 62±9 | 3.3 | 99.71 | 99.93 |
| Combined DU2 | 6.2±6.2 | 1.5±6.2 | ¡22 | NC | 1 | 41.35 | 63.19 |
| Combined DU3 | 8.1±6.2 | 1.2±6.2 | ¡24 | NC | 1.3 | 58.59 | 73.42 |